C_{3}O IN THE CARBON STAR IRC+10216

\underline{EMILY D. TENENBAUM}, ALDO J. APPONI and LUCY M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721.

We report the detection of C_{3}O in the circumstellar envelope of IRC+10216. Five rotational lines of this molecule were observed at 2 and 3 mm using the Kitt Peak 12 m telescope of the Arizona Radio Observatory (ARO). The line profiles were U-shaped, indicating that the molecule is present in the outer shell with a source size ≥ 81.6”. The derived C_{3}O column density of 1x10^{12} cm^{−2} is two orders of magnitude larger than the value predicted by the most recent chemical model of IRC+10216 (Millar, Herbst & Bettens 2000). C_{3}O is the sixth oxygen-bearing molecule that has been identified in IRC+10216, and our observations represent the second astronomical detection of C_{3}O.